A wooden box 2m long 68cm wide and 30cm high is packed with plastic blocks 10 cm long 8 cm broad and 5 cm wide.how many blocks are there in the box
1
Expert's answer
2013-03-12T13:47:25-0400
Calculate the volume of the wood box:
V1 = 200*30*68=408000 cm^3
Volume of the plastic box:
V2 = 10*5*8=400 cm^3
Denote the number of blocks as n
n=(V1)/(V2) = 1020
Answer: in the wood box 1020 blocks